OpenStudy (anonymous):

Jesse played 2 days of golf. On the second day, he got a score of 6 below par. or -6. His total score for the two days was 0 above par. or 0. Define a variable. Then write and solve an equation to find the score Jesse got on the first day. Show your work.. PLEASE HELP!!! Will Give medal and Fan

OpenStudy (anonymous):

If you played 2 game, one game you got -6 and ended with an average of 0, what does the other value have to equal

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

okay, good

OpenStudy (anonymous):

So how would you write that in an equation?

OpenStudy (anonymous):

You had to solve for 6 so lets call that "X", ill give you a hint, one part could be (-6+x) "----"=0

OpenStudy (anonymous):

if you want an average of two scores, what do you divide by?

OpenStudy (anonymous):

-6+6=0?

OpenStudy (anonymous):

yes but that doesnt account for taking an average of the two scores.

OpenStudy (anonymous):

So your saying?...

OpenStudy (anonymous):

Would it be. -6+6=0?

OpenStudy (anonymous):

if you want an average of two score, you divide by 2.

OpenStudy (anonymous):

so you would identify variable X to represent the score on the first day making the equation (-6+x)/2=0